# Advances in Tabulating Carmichael Numbers

### v1

@@ 8. Implementation, statistics, and conclusion
Here are the counts of Carmichael numbers constructed in the large preproduct case, broken down by $d$ [^1] A correction computation is being run for $d=7$. This line might not be correct as of submission. It is also possible that the correction produces no Carmichael numbers.:
$\begin{array}[]{|l|r|}\hline\cr d&C_{d}(x)\\ \hline\cr 4&100164\\ \hline\cr 5&1026622\\ \hline\cr 6&5776583\\ \hline\cr 7&14025941\\ \hline\cr 8&16008679\\ \hline\cr 9&8939992\\ \hline\cr 10&2347828\\ \hline\cr 11&262818\\ \hline\cr 12&10018\\ \hline\cr 13&55\\ \hline\cr\end{array}$
Throughout the paper, we have indicated several open questions. We intend to return to this problem and have already given thought to improved preproduct generation.

### v2

@@ 8.2. Large Preproduct Information
Here are the counts of Carmichael numbers constructed in the large preproduct case, broken down by $d$:
$\begin{array}[]{|c|c|c|c|c|}\hline\cr 4&5&6&7&8\\ \hline\cr 100164&1024718&5769410&14017455&16005059\\ \hline\cr\hline\cr 9&10&11&12&13\\ \hline\cr 8939434&2347828&262818&10018&55\\ \hline\cr\end{array}$
The following table gives timing information for the $d=6$ tabulation. Method $1$ is what we propose in Subsection 5.3 which uses the simplest example of the divisors in residue class results. Method $2$ is what is described in Subsection 5.2 and balances sieving against naive trial division. All timings are given in seconds run on a single core of an AMD Ryzen 9 3900X.
$\begin{array}[]{|c|c|c|c|c|c|c|}\hline\cr B&\mbox{Method 1 (SW)}&\mbox{Method 2 (Pinch)}\\ \hline\cr 10^{14}&25&25\\ \hline\cr 10^{15}&192&189\\ \hline\cr 10^{16}&1324&1306\\ \hline\cr 10^{17}&8752&8865\\ \hline\cr 10^{18}&55631&56072\\ \hline\cr 10^{19}&361983&364816\\ \hline\cr\end{array}$
The improvements are very modest. The reason is that the proportion of cases for which the new method gives an exponential speedup are relatively rare. We know that the new method outperforms Pinch’s at asymptotic scales, but unfortunately the scales at which computations are performed are not asymptotic. One observation that gives hope is that the timings of Method 1 got better as the upper bound increased. It is not too hard to pick out isolated examples where the improve algorithm is demonstrably better. For example, consider $P=11756813235$. The timings measured were $9$ microseconds for Method 1, and $17$ microseconds for Method 2.
8.3. Future Work
We intend to continue to work on this problem. Our guidance is that completing a preproduct is done on average in polynomial time; thus the primary cost of tabulation was constructing the preproducts. It would be better if the cost to enter the inner loop was insignificant compared to the work done in the inner loop. We believe there are faster ways of constructing preproducts and that we may consider fewer preproducts. Our goal is to tabulate all Carmichael numbers less than $10^{23}$.